OpenStudy (anonymous):

linear equation -5+7b=8b-6

OpenStudy (blurbendy):

-5 + 7b = 8b - 6 7b = 8b - 1 -b = - 1 b = 1

OpenStudy (anonymous):

ty :)

OpenStudy (blurbendy):

no prob.